Project

General

Profile

Assembly calendar: http://lariat.fnal.gov/internal/atwork.html

Cryostat

  • Phase separator (started: Nov. 14. concluded:?? )
    • Remove original insulation
    • Cut out old phase separator
    • Weld in new, larger phase separator
    • Leak/pressure check (Completed week of December 7)
    • Re-insulate piping near phase separator (Week of December 14... eventually week of Feb. 11)

Light Collection System

  • Define and implement grounding scheme for PMTs (with help from experts: Linda + Carl/Dean, in progress: started week of December 7)
    • Add non-conductive patchpanels for HV lines (Completed December 9)
    • Order new 6 or 8 pin FT (for HV) Andrzej at Manchester (Ordered, estimated arrival @ FNAL the week of February 8th)
  • Define if need screen (grid to ground) in front of the PMTs (Week of January 4)
  • Test and characterization (prior to installation):
    • Hamamatsu 3" PMT: new PMT (but same type as the one in Run1), new base & socket (change polarity from NEG to POS), new cables (signal and HV), new FT.
      • Test of electrical continuity and functioning (check for undershot). (Completed Dec. 10)
      • Test in LN2 with LED
    • ETL 2" PMT: same PMT as in Run1, same base but with modification for change polarity (from NEG to PS), new cables (signal and HV), new FT.
      • Test of electrical continuity and functioning (check for undershot) (Completed December 15)
      • Test in LN2 with LED (Week of January 18)
    • Hamamatsu SiPM (#1 and 2): same as in Run 1
    • SensL 3xSiPM new board
      • Tests and characterization performed during Summer by Dante T (WH, 14th floor) (Completed October 2015)
      • Ultimately unable to include in system for Run II.
    • (Tentative) Add new SiPM with inductive R/O from TPC wires (to be defined)
    • Add new VUV SiPM (Estimated start: Week of January 11, 2016) Finished Feb 3
  • TPB coating: one PMT (ETL) and one SiPM (HMMTS #1) should be coated with TPB+polystyrene film (toluene mixture) (Estimated: week of January 11, 2016) Finished Feb 4
  • Cabling (in situ): 2 PMTs, 3 SiPM boards - Inner Flange - Walter J. (Estimated start: week of January 18) Finished Feb 4
  • Cabling (in situ): 2 PMTs, 3 SiPM boards - Outer Flange - Walter J. (Estimated start: week of February 8)
  • DAQ & trigger modifications
    • save multiple flashes during drift time (Bill B., can start any time)
    • Implement trigger on PMT-only signal, above variable threshold (Bill B., can start any time)

TPC - readout tests

  • Modify D2S cards to alleviate analog noise, add termination on differential input lines (Completed week of Dec. 30)
  • Reinstall and test D2S cards (Completed week of Jan. 4)
  • Construct sheet metal temporary closure plate and get safety approval for operations with low HV. (Completed week of Nov. 30)
  • Channel mapping testing (checking cold cable connections) (Completed week of Jan. 4)
  • Test of the bias voltage card: Signal gen. on the wire planes, look for induced voltage on readout of appropriate plane. (Estimated: week of Jan. 4)
  • Modify bias voltage card for shield plane (Completed week of Nov. 30)-
  • Modify WRD cards to fix 200 MHz noise problem (Dean)
  • Cold electronics (on-going test - add detail)
  • Check position of TPC inside cryostat - move closer to excluder? (Estimated week of Jan. 25, before closing flanges)
  • Close up flange (Estimated week of Jan. 25)

Cosmic Tower

  • Alignment after re-installation? Good
  • Check counter's operating voltage after re-installation (needed - it'll take couple of hours) Good

Muon Range stack

  • Power up, check current draw, and check noise rate is unchanged.
  • Upgrade
    • Assemble X scintillator paddles with new small PMTs (In progress, Tapasi G. and Pip H.)
    • Install X paddles
    • Power X paddles
    • Cable X paddles to readout
    • Incorporate X and Y paddles into trigger logic.

TOF system:

  • (Jan 25-27) Stack the two paddles closely, trigger on coincidences, and measure offset. Flip stack and repeat. Average.
    • Save data for re-measurement when more sophisticated templating or CFD-like time measurement methods become available.
  • Extend time window for slow protons (Time needed: few hours, can happen any time)
  • Possibly change PMTs to reduce difference from upstream to downstream counter (low priority, can happen any time)

Beamline

  • Survey TPC in place.
  • Close cryostat flanges. (Week of Jan. 25?)
    • Remove hydraulic table
    • Reposition collimator
    • Reposition MWPCs, TOF, aerogel, Beam halo, cosmic paddles
  • Survey everything, especially that which couldn't be surveyed while TPC was open.
    • Re-cable beamline instruments
    • Test for functionality (Week of Feb. 8)

Beam returns February 5 (originally thought to be Feb. 12!)